Pi High school ordered 40 science books. The next week, the school ordered 30 algebra books. The bill for the first order was $360 greater than the bill for the second order. The two bills together totaled $3960. Find the price of the algebra book.
Answer by macston(5194)

A=bill for algebra books; S=bill for science books=A+$360
S+A=$3960 Substitute for S
A+$360+A=$3960 Subtract $360 from each side.
2A=$3600 Divide each side by 2.
A=$1800 The order for 30 algebra books cost $1800.
Individual book=A/30
individual book=$1800/30=$60 ANSWER: Each algebra book cost $60.